the doctor orders strepto mycin 1.75 mg/lb IM every 12 hrs. Available is Streptomycin 0.35 g/2.3 mL. How many mL will you administer a day to a 59 Kg patient?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug administer per day.

       Formula

          Volume(ml) = Desired dose(D) / Dose in hand(H) * Quantity(Q)

Here

        Ordered dose = 1.75 mg/lb every 12 hrs.

        Weight of the patient = 59 kg.

   Convert kg to lb for calculation

              1 kg = 2.205 lb.

   So

      Weight of the patient in lb = 59 * 2.205

                                               = 130.09 lb.

        Desired dose(D) for this patient = 1.75 * 130 .09

                                                            = 227.66 mg.(every 12hrs)

                       Dose in hand (H) = 0.35 gm.

          Convert gm to mg for calculation

                   1 gm = 1000 mg.

       So

               Dose in hand (H) = 0.35 * 1000

                                           = 350 mg.

                      Quantity(Q) = 2.3 ml

As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 227.66 / 350 * 2.3

                                                        = 0.65 * 2.3

                                                        = 1.49

                                                        = 1.5 ml every 12 hrs.

          Every 12 hrs is 2 times a day

              So

                 Volume(ml) of drug required per day = 1.5 * 2

                                                                            = 3 ml per day.